OpenStudy (anonymous):

Please solve log4 (y – 9) + log4 3 = log4 81

jhonyy9 (jhonyy9):

property of logarithm log a + log b =log (a*b)

jhonyy9 (jhonyy9):

so ((y-9)3=81 3y-27=81 3y=108 y=108/3 y=36